# EasyOcc: 3D Pseudo-Label Supervision for Fully Self-Supervised Semantic Occupancy Prediction Models

### v3

@@ V-C Main Results: Camera Mask
In this section, we evaluate the performance of seven baseline models against the modified versions of the three selected baseline models and EasyOcc. We compare the models across four key evaluation categories: inference time (FPS), Intersection over Union (IoU), mean IoU (mIoU), and class-wise IoU for each semantic category. Results are provided in Table III.
@@ V-C1 Inference Time
GaussianFlowOcc achieves the highest inference speed at 10.2 FPS, attributed to its use of induced attention, which significantly reduces computational overhead. SelfOcc ranks second with 7.4 FPS, benefiting from the lack of a contracted coordinate system. The inclusion of pseudo-loss has no effect on inference time, as it influences only the training phase. EasyOcc matches the inference speed of both OccNeRF and GaussianOcc, all of which operate at 5.4 FPS.
V-C2 Intersection over Union (IoU)
@@
AutoOcc secures the highest performance with an IoU of 83.01, significantly surpassing all other models. GaussianFlowOcc ranks second, achieving an IoU of 46.91. Notably, the addition of pseudo-loss led to a marked decline in IoU, for example, it reduces SelfOcc’s score from 44.05 to 34.50, a 22% drop. This performance drop may be attributed to object duplication, as the model lacks the ability to reason about occluded regions. Consequently, it tends to predict occupancy beyond visible surfaces, resulting in an overly dense scene representation. This behavior is visualized in the qualitative analysis presented in Subsection V-F, and further substantiated by evaluations without the camera mask, assessing every voxel in the space, which shall be discussed in Subsection V-D.
V-C3 Mean Intersection over Union (mIoU)
@@
A notable performance gap exists between the original voxel-based models (SelfOcc, OccNeRF, and GaussianOcc) and the Gaussian-based models: GaussTR, TT-OccCamera, GaussianFlowOcc, and AutoOcc. AutoOcc once again leads with an mIoU of 20.92, largely due to the integration of a VLM [InternVL] and a VFM [fastsam] during inference, which provides high-quality semantic predictions.
@@
The incorporation of pseudo-loss enables substantial improvements to existing models, surpassing even the Gaussian-based methods: GaussTR and TT-OccCamera. When pseudo-loss is applied to OccNeRF, it achieves a 15% improvement over GaussTR and a 45% gain compared to the original OccNeRF model. Similar performance boosts are observed for both SelfOcc and GaussianOcc. Notably, SelfOcc gains the ability to predict previously unsupported classes, such as construction vehicles, thanks to semantic supervision from Grounded-SAM.
Our model, EasyOcc, reaches an mIoU of 13.86, placing it on par with the other three models that utilize rendering-based supervision along with pseudo-loss.
TABLE IV: Model Comparison evaluated without the Camera Mask: Models are retrained from Table III without the camera mask. FPS denotes frames per second, representing the time taken to process a single sample. IoU and mIoU refer to Intersection over Union and mean Intersection over Union, respectively. The best result in each category is highlighted in bold, while the second-best is underlined. * OccNeRF implements 2D semantic loss, whereas GaussianOcc does not.
@@
Model FPS IoU mIoU $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [selfocc] 7.4 6.96 2.78 0.43 0.45 7.45 5.58 0.00 0.07 1.96 0.00 0.00 5.45 5.65 6.55 4.73 2.84 3.33 OccNeRF [occnerf] 5.4 5.95 2.57 0.03 0.16 2.42 1.28 0.73 0.03 0.37 0.13 0.45 0.92 21.08 5.11 2.84 3.88 4.33 GaussianOcc [gaussianocc] 5.4 5.33 5.90 0.59 3.27 11.94 7.99 0.49 1.43 5.04 5.19 0.32 7.28 24.34 13.05 10.55 1.42 7.47 GaussTR [gausstr] 0.3 11.58 4.37 1.49 1.96 12.22 9.13 1.67 6.58 3.30 1.09 1.56 8.62 9.33 3.72 2.12 5.00 6.43 SelfOcc [selfocc] + Ours 7.4 18.22 7.91 0.80 2.65 10.33 15.30 1.45 3.25 10.37 4.03 0.56 8.99 27.27 15.32 12.06 6.47 7.78 OccNeRF [occnerf] + Ours * 5.4 16.88 8.21 0.84 5.16 11.62 14.71 0.56 4.94 11.72 7.04 0.66 9.34 27.37 16.22 12.37 6.71 10.24 GaussianOcc [gaussianocc] + Ours * 5.4 17.13 7.77 0.80 2.98 10.47 14.81 0.95 4.34 11.32 4.63 0.57 8.74 27.60 15.88 12.03 6.84 10.22 EasyOcc (Ours) RN-101 5.4 16.98 7.71 0.86 3.22 9.57 14.46 1.14 4.31 11.07 5.08 0.74 8.32 27.48 15.89 11.83 6.83 10.28
@@ V-C4 IoU per Semantic Class
AutoOcc, GaussianFlowOcc, and GaussTR achieve the highest performance across all semantic classes. AutoOcc leads in 6 out of 15 classes, particularly excelling in smaller object categories such as bicycle, motorcycle, and traffic cone, likely due to the use of foundation models during inference. GaussianFlowOcc ranks first in 4 classes, primarily large-scale categories such as drivable surface, sidewalk, and terrain, which may be attributed to the strong emphasis on temporal information during training. GaussTR attains top performance in 5 classes, notably dynamic objects such as bus and car, which is an interesting outcome given the absence of a dedicated flow module or temporal modeling, which GaussianFlowOcc and AutoOcc implement.
@@
None of the voxel-based models, SelfOcc, OccNeRF, GaussianOcc, EasyOcc, or their pseudo-loss-enhanced variants achieves top performance in any class. Nevertheless, with the integration of pseudo-loss, these models deliver competitive results, securing second place in several classes. Significant improvements in IoU are observed across nearly all categories. For GaussianOcc, IoU increases in 14 out of 15 classes, with the barrier class as the sole exception. Similar trends are seen in SelfOcc and OccNeRF. Notably, SelfOcc demonstrates a dramatic 624% increase in IoU for the pedestrian class, emphasizing the importance of 3D labels for accurately detecting vulnerable object categories. EasyOcc exhibits strong detection capabilities, achieving performance comparable to that of the modified models enhanced with pseudo-loss.
V-D Main Results: No Camera Mask
In this section, we re-examine the previously reported results, this time excluding the use of the camera mask, as motivated in earlier Subsection IV-D. The models are evaluated in the updated results Table IV based solely on IoU, mIoU, and per-class IoU. FPS is omitted as it remains the same as the previous. TT-OccCamera, GaussianFlowOcc, and AutoOcc are excluded from this analysis due to the lack of publicly available implementations.
V-D1 Intersection over Union (IoU)
Compared to evaluation with the camera mask, we see a stark drop in performance, most notably for the original models, where GaussTR observes a 74% decrease. EasyOcc observes much less of a decrease of 56%, achieving a higher overall IoU of 16.98 compared to GaussTR’s 11.58. All models incorporating pseudo-loss experience massive increases in IoU, with SelfOcc achieving the highest IoU of 18.22, a 162% increase over the original SelfOcc’s IoU of 6.96. This is in contrast to the camera mask models, where pseudo-loss worsened IoU metrics, displaying that our labels do indeed enable a better scene representation.
V-D2 Mean Intersection over Union (mIoU)
The trend continues into mIoU, where again the original models suffer much greater on average than our models, when compared to evaluation with the camera mask. One outlier is GaussianOcc, which observes a 41% decrease, whereas GaussianOcc with pseudo-loss observes a 44% decrease. However, our models still achieve the best performance, with OccNeRF including our 3D pseudo-labels achieving 8.21 mIoU. EasyOcc achieves an mIoU of 7.71, despite using many fewer loss computation techniques than OccNeRF.
We note that both SelfOcc and OccNeRF take the biggest hit in performance when evaluating on the entire grid. This likely originates from their overprediction and volume rendering, which causes object duplication, heavily penalizing the mIoU metric regarding false positives (FP), as seen in Equation (11). This is more evident when examining IoU per semantic class, and is further evident in the qualitative analysis Subsection V-F1.
V-D3 IoU per Semantic Class
Here, we examine a clear distinction between the models that utilize the 3D pseudo-labels and the ones that do not. SelfOcc, OccNeRF, and GaussianOcc fail to achieve top performance across any class, whereas GaussTR successfully achieved top performance on 5 out of 16 total classes. However, the remaining top-performing models in all other classes use 3D pseudo-labels. For example, when applying our 3D pseudo-labels to OccNeRF, the model outperforms GaussTR in 10 out of 16 semantic classes and achieves the highest performance on six classes across all evaluated models.
Looking further into specific classes, we find that 3D pseudo-labels enable much better detection on all classes. For pedestrian detection, GaussianOcc achieved 5.04 mIoU, and this is increased to 11.32 mIoU when 3D pseudo-labels are used for supervision. For the original models, some classes do achieve respectable performance. For example, GaussianOcc achieves an mIoU of 24.34 for drivable surface, and this increases to just 27.60 mIoU when 3D pseudo-labels are implemented. A similar observation is seen for OccNeRF. For SelfOcc, all semantic classes show improved performance with the integration of our 3D pseudo-labels. In the case of OccNeRF, all classes improve except for construction vehicles, while for GaussianOcc, all classes improve except for bicycle, bus, and traffic cone. Notably, the decline in mIoU for these few classes is marginal.
For EasyOcc, we examine performance comparable with that of the models incorporating pseudo-loss, achieving top performance for one category only, vegetation, due to the dominance of OccNeRF with pseudo-labels.
@@ V-E Ablation Study: EasyOcc
In this section, we present a series of ablation studies on the EasyOcc model to assess the impact of key design choices. Models are trained for 12 epochs, balancing thorough evaluation with practical constraints on energy consumption and training time. First, in Subsubsection V-E1, we examine the effect of varying the image encoder and input image resolution. This is followed by an analysis of temporal sample aggregation in Subsubsection V-E2. Next, in Subsubsection V-E3, we evaluate the sensitivity of pseudo-loss to changes in the weighting constant $\lambda$. In Subsubsection V-E4, we investigate the effect of selectively removing components of the pseudo-loss. Finally, in Subsubsection V-E5, we explore the performance of Lovász-Softmax loss in more detail. All evaluation is performed with the camera mask.
@@ V-E1 Image Encoder and Image Size
In Table V, we investigate the effect of varying the image encoder and input image resolution on the mIoU metric, total model parameters, and inference speed (FPS).
TABLE V: Ablation: Image Encoder and Input Image Size: Models are trained for 12 epochs to ensure reasonable training time. The best-performing model in each category for each input resolution is highlighted in bold.
@@
While shallower models typically exhibit higher inference speeds, this trend is less prominent in our results due to the inference bottleneck introduced by grid sampling, inherited from the contracted coordinate system used in OccNeRF and GaussianOcc. Thus, for a marginal trade-off in inference speed, deeper models like RN-101 or RN-152 may be preferable. However, deeper models also come with a substantial increase in parameter count—transitioning from RN-18 to RN-152 results in a 423% increase in parameters. Therefore, selecting the encoder requires balancing performance, inference speed, and memory constraints.
@@ V-E2 Aggregation of Temporal Samples
Figure 7: Ablation: Temporal Samples: Models are trained for 12 epochs to maintain manageable training time.
The figure exhibits a logarithmic curve, seen previously when comparing Occ3D ground-truth labels to the 3D pseudo-labels, with the results clearly showing that increasing the number of temporal samples improves the mIoU metric. This improvement is attributed to scene densification, especially in regions farther from the camera’s field of view. Notably, we achieve an mIoU of 13.18 with both 10 and 13 temporal samples, further indicating a saturation point beyond which additional samples yield diminishing returns in the mIoU metric.
@@ V-E3 Variation of $\lambda$
In Table VI, we experiment with the $\lambda$ constant in Equation (5) to balance the contributions of the cross-entropy and geometry loss components. The results show that a $\lambda$ value of 0.1 yields the best mIoU performance. A value of 1 performs slightly worse, trailing by 0.22 points, while a value of 0.01 results in a significant drop of 0.80 points. These findings indicate that geometry loss plays a meaningful role, but balancing the contributions of both cross-entropy and geometry losses is crucial. With $\lambda=0.1$, the magnitudes of the cross-entropy and geometry losses are approximately equal during testing, underscoring the importance of careful weighting for optimal performance.
TABLE VI: Ablation: Variation of $\lambda$ in Pseudo-Loss: Models are trained for 12 epochs to ensure practical training time. The best-performing configuration is highlighted in bold.
@@ V-E4 Choice of Losses
In Table VII, we ablate the pseudo-loss in Equation (5) by removing individual components, such as the cross-entropy term, and elements of the geometry loss in Equation (6), to assess the contribution of each component to the model’s learning.
TABLE VII: Ablation: Dropping Losses in Pseudo-Loss: Models are trained for 12 epochs to maintain feasible training time. The best-performing configuration is highlighted in bold.
@@
13.18
13.29
12.21
12.54
11.02
8.15
8.71
Overall, the results show that removing components of the loss function reduces the model’s learning capability. For example, excluding the cross-entropy loss decreases mIoU by 0.97 points. Similar drops occur when individual components of the geometry loss are removed. Interestingly, excluding the Lovász-Softmax loss results in a slight mIoU improvement of 0.11, which is unexpected since it is designed to optimize the IoU metric. We hypothesize this anomaly arises from including the empty class index (17) in the Lovász loss computation. To validate this, we retrain the model excluding this index, as detailed in the following ablation.

### v4

@@ V-C MAIN RESULTS: MIOU
In this section, we evaluate the performance of the three selected baseline models and EasyOcc in Table III. We compare the models across four key evaluation categories: inference time (FPS), Intersection over Union (IoU), mean IoU (mIoU), and class-wise IoU for each semantic category. We provide further models for reference in the table.
@@ V-C1 INFERENCE TIME
GaussianFlowOcc achieves the highest inference speed at 10.2 FPS, attributed to its use of induced attention, which significantly reduces computational overhead. SelfOcc ranks second with 7.4 FPS, benefiting from the lack of a contracted coordinate system. The inclusion of pseudo-loss has no effect on inference time, as it influences only the training phase. EasyOcc matches the inference speed of both OccNeRF and GaussianOcc, all of which operate at 5.4 FPS. Both GaussTR and TT-Occ exhibit poor FPS due to the deployment of foundation models at inference time.
V-C2 INTERSECTION OVER UNION
@@
AutoOcc secures the highest performance with an IoU of 83.01, significantly surpassing all other models. GaussianFlowOcc ranks second, achieving an IoU of 46.9. Notably, the addition of pseudo-loss led to a marked decline in IoU; for example, it reduces SelfOcc’s score from 44.1 to 34.5, a 22% drop. This performance drop is attributed to object duplication, as the model lacks the ability to reason about occluded regions. Consequently, it tends to predict occupancy beyond visible surfaces, resulting in an overly dense scene representation, which is further evidenced in the evaluation of RayIoU in Subsection V-D. This is visualized in the qualitative analysis presented in Subsection V-F.
V-C3 MEAN INTERSECTION OVER UNION
@@
A notable performance gap exists between the original voxel-based models (SelfOcc, OccNeRF, and GaussianOcc) and the Gaussian-based models: GaussTR, TT-Occ, GaussianFlowOcc, and AutoOcc. AutoOcc once again leads with an mIoU of 20.9, largely due to the integration of a VLM [8] and a VFM [60] during inference, which provides high-quality semantic estimations.
@@
With the sole usage of Easy3D-Labels, EasyOcc reaches an mIoU of 15.7, surpassing even the Gaussian-based methods GaussTR. The employment of pseudo-loss allows OccNeRF to achieve a 15% improvement over GaussTR and a 45% gain compared to the original OccNeRF model. Similar performance boosts are observed for both SelfOcc and GaussianOcc. Notably, SelfOcc gains the ability to predict previously unsupported classes, such as construction vehicles, thanks to semantic supervision from Grounded-SAM. Furthermore, despite the lack of usage of rendering for loss, EasyOcc achieves results on par with the other three models, which use Easy3D-Labels, displaying the strength of our labels.
TABLE IV: RayIoU evaluated on the Occ3D-nuScenes [45] dataset: * OccNeRF implements 2D semantic loss, whereas GaussianOcc does not. The best-performer is highlighted in bold, for each model compared to the same model trained with our labels integrated.
@@
Method RayIoU RayIoU@1 RayIoU@2 RayIoU@4 $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [17] 9.6 6.7 9.6 12.5 0.7 1.0 15.9 16.3 0.0 0.9 5.0 0.0 0.0 20.8 48.7 13.4 17.2 15.9 8.0 + Ours 14.5 10.9 14.6 17.9 3.1 4.9 24.1 29.7 3.4 5.6 23.3 13.7 1.0 27.2 49.1 17.2 13.7 15.2 17.8 OccNeRF [55] 10.4 6.9 10.3 14.1 2.4 2.2 28.4 21.5 4.4 1.0 5.6 7.5 0.6 16.8 38.8 11.8 10.3 13.0 10.0 + Ours 15.5 11.7 15.6 19.2 3.2 6.5 27.2 30.0 3.2 6.5 23.5 16.2 1.9 29.5 50.2 18.0 14.6 15.1 20.2 GaussianOcc [9] 11.9 8.7 11.9 15.0 2.6 7.0 24.2 18.3 2.0 2.6 12.1 13.3 0.7 23.6 41.7 15.9 14.0 11.5 12.8 + Ours 14.3 10.7 14.4 17.8 2.9 3.8 23.4 29.4 3.6 6.2 21.8 12.3 0.8 25.7 48.9 17.4 13.5 15.0 20.0 EasyOcc (Ours) 14.6 10.9 14.7 18.2 3.1 4.9 22.7 28.7 4.8 6.6 21.4 13.2 2.4 27.4 49.8 17.2 13.3 14.8 19.9
TABLE V: RayIoU evaluated on the OpenOccv2 [46] dataset: * OccNeRF implements 2D semantic loss, whereas GaussianOcc does not. The best-performer is highlighted in bold, for each model compared to the same model trained with our labels integrated.
Method RayIoU RayIoU@1 RayIoU@2 RayIoU@4 $\blacksquare$ barrier $\blacksquare$ bicycle $\blacksquare$ bus $\blacksquare$ car $\blacksquare$ const. veh. $\blacksquare$ motorcycle $\blacksquare$ pedestrian $\blacksquare$ traffic cone $\blacksquare$ trailer $\blacksquare$ truck $\blacksquare$ drive. surf. $\blacksquare$ sidewalk $\blacksquare$ terrain $\blacksquare$ manmade $\blacksquare$ vegetation SelfOcc [17] 9.1 5.8 9.1 12.3 0.6 0.9 14.4 15.3 0.0 0.8 4.8 0.0 0.0 20.1 39.0 11.8 13.6 15.7 7.3 + Ours 15.1 11.2 15.3 18.7 2.9 4.3 23.2 28.5 3.5 5.1 22.7 12.4 1.3 26.6 45.4 16.3 18.0 16.2 19.2 OccNeRF [55] 11.4 7.9 11.3 15.0 2.1 1.8 29.4 21.6 4.3 0.9 4.9 6.2 0.6 17.3 41.8 13.8 13.8 12.7 9.6 + Ours 15.9 12.0 16.1 19.6 2.9 5.3 25.8 28.3 3.4 5.9 22.5 14.3 2.0 28.9 46.5 17.4 18.5 15.4 21.0 GaussianOcc [9] 11.7 8.5 11.8 14.8 2.4 5.8 22.8 17.7 2.0 2.2 11.0 11.1 0.6 23.3 37.8 14.9 13.7 11.1 12.2 + Ours 14.9 11.1 15.1 18.5 2.6 3.1 22.4 28.0 4.0 5.8 20.8 10.7 1.1 24.9 47.0 17.3 18.1 15.2 20.7 EasyOcc (Ours) 15.1 11.3 15.4 18.7 2.9 4.3 21.6 27.3 5.0 6.0 20.6 11.4 3.1 26.5 46.7 17.1 18.2 15.0 20.7
@@ V-C4 IOU PER SEMANTIC CLASS
For single-class mIoU, AutoOcc leads in 6 out of 15 classes, particularly smaller objects such as bicycle, motorcycle, and traffic cone, likely due to the use of foundation models during inference. GaussianFlowOcc ranks first in 3 classes, mainly large-scale categories like drivable surface, sidewalk, and terrain, likely benefiting from strong temporal modeling. GaussTR achieves top performance in 5 classes, especially dynamic objects such as buses and cars, despite not using a dedicated flow module or temporal modeling.
@@
With the integration of Easy3D-Labels into SelfOcc, OccNeRF, and GaussianOcc, results increase significantly across nearly all categories. For GaussianOcc, IoU increases in 14 out of 15 classes, with the barrier class as the sole exception. Similar trends are seen in SelfOcc and OccNeRF. We observe that performance generally improves for dynamic classes, while some static classes, particularly large-area categories such as manmade, do not show the same gains, potentially due to overprediction aiding their segmentation in models without our labels. Notably, SelfOcc demonstrates a dramatic 624% increase in IoU for the pedestrian class, emphasizing the importance of 3D labels for accurately detecting vulnerable object categories. EasyOcc exhibits strong detection capabilities, achieving performance comparable to that of the modified models enhanced with pseudo-loss.
V-D MAIN RESULTS: RAYIOU
In this section, we analyse models on the RayIoU metric across varying distance thresholds, along with class-wise RayIoU in Tables IV and V. As discussed previously, this metric penalises overprediction, which can otherwise inflate mIoU without accurately reflecting overall scene understanding. Additionally, evaluation across two datasets enables a more robust and consistent analysis.
Across both tables, we observe that training with Easy3D-Labels results in a substantial improvement in RayIoU for all distance thresholds. For example, OccNeRF shows a 49% and 39% increase on the Occ3D and OpenOccv2 datasets, respectively. Similar to the mIoU results, OccNeRF achieves the highest overall performance, reaching RayIoU scores of 15.5 and 15.9. Evaluating across different distance thresholds also highlights improved depth accuracy, with RayIoU@1 for SelfOcc increasing by over 93% on the OpenOccv2 dataset. Furthermore, these results emphasise the overprediction issue present in the original models, where IoU scores were artificially inflated, and for which RayIoU penalizes.
For class-wise RayIoU, Easy3D-Labels again enable more accurate detection, consistent with trends observed in the mIoU results, particularly for dynamic objects. Vulnerable classes, such as pedestrians, show significant improvements; for example, OccNeRF on Occ3D achieves an increase of over 300%, largely due to the accuracy provided by Metric3Dv2 [14]. Additionally, SelfOcc on OpenOccv2 shows improvements across all classes, with notable gains for car, motorcycle, and pedestrian. Our model, EasyOcc, demonstrates strong and competitive performance compared to SelfOcc trained with Easy3D-Labels, despite not relying on rendering-based losses.
@@ V-E ABLATION STUDY: EASYOCC
In this section, we present a series of ablation studies on the EasyOcc model to assess the impact of key design choices. We ablate the image encoder, temporal samples, and pseudo-loss, among others. Models are trained for 12 epochs for computational efficiency.
@@ V-E1 IMAGE ENCODER AND IMAGE SIZE
In Table VI, we investigate the effect of varying the image encoder and input image resolution on the mIoU metric, total model parameters, and inference speed (FPS).
TABLE VI: Image encoder and input image size ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
@@ V-E2 AGGREGATION OF TEMPORAL SAMPLES
Figure 6: Temporal sample ablation.
The figure exhibits a logarithmic curve, seen previously when comparing Occ3D ground-truth labels to the 3D pseudo-labels, with the results clearly showing that increasing the number of temporal samples improves the mIoU metric. This improvement is attributed to scene densification, especially in regions farther from the camera’s field of view. Notably, we achieve an mIoU of 14.9 with both 10 and 13 temporal samples, further indicating a saturation point beyond which additional samples yield diminishing returns in the mIoU metric.
@@ V-E3 VARIATION OF $\lambda$
In Table VII, we experiment with the $\lambda$ constant in Equation (5) to balance the contributions of the cross-entropy and geometry loss components. The results show that a $\lambda$ value of 0.1 yields the best mIoU performance. A value of 1 performs slightly worse, trailing by 0.2 points, while a value of 0.01 results in a significant drop of 0.7 points. These findings indicate that geometry loss plays a meaningful role, but balancing the contributions of both cross-entropy and geometry losses is crucial. With $\lambda=0.1$, the magnitudes of the cross-entropy and geometry losses are approximately equal during testing, underscoring the importance of careful weighting for optimal performance.
TABLE VII: Pseudo-loss $\lambda$ ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
@@ V-E4 CHOICE OF LOSSES
In Table VIII, we ablate the pseudo-loss in Equation (5) by removing individual loss components in Equation (6), to assess the contribution of each component to the model’s learning.
TABLE VIII: Pseudo-loss ablation: A grey row color denotes the choice for the final model. The best-performing model in each category for each input resolution is highlighted in bold.
@@
14.9
15.1
13.8
14.2
12.5
9.2
9.9
Overall, the results show that removing components of the loss function reduces the model’s learning capability. For example, excluding the cross-entropy loss decreases mIoU by 1.1 points. Similar drops occur when individual components of the geometry loss are removed. Interestingly, excluding the Lovász-Softmax loss results in a slight mIoU improvement of 0.2, which is unexpected since it is designed to optimize the IoU metric. We hypothesize that this anomaly arises from including the empty class index in the Lovász loss computation. To validate this, we retrain the model excluding this index, as detailed in the following ablation.